# SubIQ: Inverse Soft-Q Learning for Offline Imitation with Suboptimal Demonstrations

### v1

@@ Abstract
We consider offline imitation learning (IL), which aims to mimic the expert’s behavior from its demonstration without further interaction with the environment. One of the main challenges in offline IL is dealing with the limited support of expert demonstrations that cover only a small fraction of the state-action spaces. In this work, we consider offline IL, where expert demonstrations are limited but complemented by a larger set of sub-optimal demonstrations of lower expertise levels. Most of the existing offline IL methods developed for this setting are based on behavior cloning or distribution matching, where the aim is to match the occupancy distribution of the imitation policy with that of the expert policy. Such an approach often suffers from over-fitting, as expert demonstrations are limited to accurately represent any occupancy distribution. On the other hand, since sub-optimal sets are much larger, there is a high chance that the imitation policy is trained towards sub-optimal policies. In this paper, to address these issues, we propose a new approach based on inverse soft-Q learning, where a regularization term is added to the training objective, with the aim of aligning the learned rewards with a pre-assigned reward function that allocates higher weights to state-action pairs from expert demonstrations, and lower weights to those from lower expertise levels. On standard benchmarks, our inverse soft-Q learning significantly outperforms other offline IL baselines by a large margin.
Machine Learning, ICML
@@ 1 Introduction
Reinforcement learning (RL) has been raised as a strong framework for decision-making by its ability to learn by interacting with the environment. There are several applications in diverse domains: robotics (Kilinc & Montana, 2022; Hoang et al., 2023b), healthcare (Weng et al., 2017; Raghu et al., 2017), and structure generating (Dennis et al., 2020). However, to achieve high performance, original RL methods must balance the trade-off between exploration and exploitation. To handle this, several works utilize external support, such as expert demonstrations, to guide the learning scheme to achieve better performance, which is typically referred to as imitation learning (Ho & Ermon, 2016; Reddy et al., 2019; Garg et al., 2021; Hoang et al., 2023a).
Although having these assists, the number of interactions with the environment remains extremely large. Several works are trying to tackle the problem of offline schemes with a dataset of expert demonstrations (Wu et al., 2019a; Kumar et al., 2020; Zhang et al., 2021; Ajay et al., 2022). Starting with Behavioral Cloning (BC), it aims to maximize the likelihood of the expert’s decisions from the provided dataset. Although BC has a highly theoretical background that tends to converge towards expert behaviors, it often suffers from the distributional shift between the training dataset and the environment when there are insufficient demonstrations (Ross et al., 2011).
There are several works that aim to utilize non-expert demonstration datasets to reduce the reliance on a large number of expert demonstrations. This is a well-known approach in offline imitation learning, known as using supplementary data. However, it is obvious that directly applying BC to these larger supplementary datasets can lead to sub-optimal policies. Most prior work attempts to extract expert-like demonstrations from the supplementary dataset in order to expand the dataset available for Behavioral Cloning (Sasaki & Yamashina, 2020; Kim et al., 2021, 2022; Xu et al., 2022; Yu et al., 2023). However, these works assume expert demonstrations are present in the supplementary dataset and focus on identifying and utilizing those. The rest of the dataset is eliminated. On the other hand, we learn a Q function where the sub-optimal demonstrations also contribute to our understanding of the environment and ground truth reward function.
In this work, we aim to leverage ranked non-expert datasets to address the aforementioned issues and reduce the need for expert demonstrations. Our proposed approach is based on inverse soft Q functions, incorporating a reward regularization term. This term is designed to align learning rewards with a predetermined reward function, assigning weights to state-action observations of varying expertise levels. It is worth noting that prior works (Brown et al., 2019, 2020; Chen et al., 2021) have also attempted to minimize the required demonstrations using ranked datasets by preference-based RL. However, their approaches are only applicable to online training.
Contributions: We make the following contributions:
@@
(i) We propose sub-IQ, a novel algorithm for offline imitation learning with expert and non-expert demonstrations.
(ii) We provide key theoretical properties of the Sub-IQ objective function, which enable a scalable approach.
(iii) To mitigate the overestimation issue, a common problem in offline Q learning, we develop a conservative version of Sub-IQ and demonstrate that the new objective is not only convex within the Q-space but also guarantees to return a Q function that lower-bounds its true value.
(iv) We provide extensive experiments on two benchmark domains: Mujoco and Panda-gym, showing that our algorithms consistently outperform other existing algorithms.
@@ 1.1 Related Work
Imitation learning is recognized as a significant technique for learning from expert demonstrations. It begins with Behavioral Cloning (BC), which aims to maximize the likelihood of expert demonstrations. However, BC often underperforms in practice due to unforeseen scenarios (Ross et al., 2011). To overcome this limitation, Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016) and Adversarial Inverse Reinforcement Learning (AIRL) (Fu et al., 2017) have been developed. These methods align the occupancy distributions of the policy and the expert within the Generative Adversarial Network (GAN) framework (Goodfellow et al., 2014). Alternatively, Soft Q Imitation Learning (SQIL) (Reddy et al., 2019) bypasses the complexities of adversarial training by assigning a reward of +1 to expert demonstrations and 0 to the others, subsequently learning a value function based on these rewards.
@@
While the aforementioned imitation learning algorithms show promise, they require interaction with the environment to obtain the policy distribution, which is often impractical. ValueDICE (Kostrikov et al., 2019) introduces a novel approach for off-policy training, suitable for offline training, using Stationary Distribution Corrections (Nachum et al., 2019a, b). However, ValueDICE necessitates adversarial training between the policy and the Q function, a challenge that O-NAIL (Arenz & Neumann, 2020) successfully addresses. Recently, algorithms like PWIL (Dadashi et al., 2021) and IQ-learn (Garg et al., 2021) have utilized distribution distance, offering an alternative to adversarial training schemes. Since such approaches rely on occupancy distribution matching, a large expert dataset is often required to achieve the desired performance.
Learning from Imperfect Demonstration.
Recent works have shown that utilizing imperfect demonstrations in online training schemes can provide additional information with several advantages. T-REX (Brown et al., 2019) and D-REX (Brown et al., 2020) have shown that understanding noise-ranked demonstrations as a reference-based approach can return a better policy without requiring expert demonstrations. Moreover, there are also several works (Wu et al., 2019b; Wang et al., 2021) that utilize the GAN framework (Goodfellow et al., 2014) for sub-optimal datasets and have achieved several successes.
Meanwhile, in the offline imitation learning context, from the sub-optimal demonstrations, TRAIL (Yang et al., 2021) use sup-optimal demonstrations to learn the environment’s dynamics using a feature encoder to map the high-dimensional state-action space into a lower dimension. This approach may face challenges in complex environments where predicting dynamics accurately is difficult. Other works assume that they can attract expert-like state-action pairs from the sub-optimal demonstration set and use them for BC with importance sampling (Sasaki & Yamashina, 2020; Xu et al., 2022; Yu et al., 2023; Kim et al., 2021, 2022). However, these studies must assume that expert state-actions are included in the sub-optimal set, which is not always the case. Additionally, expert-like state-actions may be misidentified, as reward information is not available. In contrast, our approach is more general, as we do not assume that the sub-optimal set contains expert demonstrations. We also allow for the inclusion of demonstrations of various qualities. Additionally, while prior works only recover expert policies, our approach enables the recovery of both expert policies and rewards, which facilitates applications in varied situations.
@@ 2 Background
We consider a MDP define by the following tuple $\mathcal{M}=\left\langle S,A,r,P,\gamma,s_{0}\right\rangle$, where $S$ denotes the set of states, $s_{0}$ represents the initial state set, $A$ is the set of actions, $r:S\times A\rightarrow\mathbb{R}$ defines the reward function for each state-action pair, and $P:S\times A\rightarrow S$ is the transition function, i.e., $P(s^{\prime}|s,a)$ is the probability of reaching state $s^{\prime}\in S$ when action $a\in A$ is made at state $s\in S$, and $\gamma$ is a discount factor. In reinforcement learning (RL), the aim is to find a policy that maximizes the expected long-term accumulated reward
$\small\max_{\pi}~{}\left\{\mathbb{E}_{(s,a)\sim\rho_{\pi}}[r(s,a)]\right\}$
where $\rho_{\pi}$ is the occupancy measure of policy $\pi$, computed as. $\rho_{\pi}(s,a)=(1-\gamma)\pi(a|s)\sum_{t=1}^{\infty}\gamma^{t}P(s_{t}=s|\pi).$
MaxEnt RL
The MaxEnt RL framework aims to optimize a policy that maximizes an entropy regularized MDP
$\max_{\pi}~{}~{}\mathbb{E}_{(s,a)\sim\rho_{\pi}}[r(s,a)]+H(\pi)$
where $H(\pi)$ is an entropy regularizer, defined as a discounted causal entropy of the learning policy $\pi$, $H(\pi)=\mathbb{E}_{\rho_{\pi}}[-\log\pi(s,a)]$. The above MDP problem yields a unique solution $\pi^{*}$ as $\pi^{*}(s,a)=\frac{e^{Q^{*}(s,a)}}{\sum_{a^{\prime}}e^{Q^{*}(s,a^{\prime})}},$ where the soft Q-function $Q^{*}$ is a unique fixed point solution for the Bellman equations ${\mathcal{B}}_{r}[Q]=Q$, where
${\mathcal{B}}_{r}[Q](s,a)=r(s,a)$
$+\gamma\mathbb{E}_{s^{\prime}\sim P(s^{\prime}|s,a)}\left[\log\left(\sum_{a^{\prime}}e^{Q(s^{\prime},a^{\prime})}\right)\right]$ (1)
@@
In the context if IRL, the objective is to recover a reward function $r(s,a)$ from a set of expert demonstrations, ${\mathcal{D}}^{E}$. Let $\rho^{E}$ be the occupancy measure of the expert policy. The MaxEnt IRL framework proposes to recover the expert reward function by solving
@@
$-\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)])\Big{\}}$ (2)
Intuitively, we aim to find a reward function that achieves the highest difference between the expected value of the expert policy and the highest expected value among all other policies (computed through the min loop).
@@
Given a reward function $r$ and a policy $\pi$, the soft Bellman equation is defined as
${\mathcal{B}}^{\pi}_{r}[Q](s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}}[V^{\pi}(s^{\prime})]$
where $V^{\pi}(s)=\mathbb{E}_{a\sim\pi(a|s)}[Qs,a)-\log\pi(a|s)]$. The Bellman equation ${\mathcal{B}}^{\pi}_{r}[Q]=Q$ is contractive and always yields a unique Q solution (Garg et al., 2021). In IQ-learn, they further define an inverse soft-Q Bellman operator
${\mathcal{T}}^{\pi}[Q]=Q(s,a)-\gamma\mathbb{E}_{s^{\prime}}[V^{\pi}(s^{\prime})]$
@@
where $V^{\pi}(s)=\mathbb{E}_{a\sim\pi(a|s)}[Q(s,a)-\log\pi(a|s)]$. (Garg et al., 2021) show that for any reward function $r(a,s)$, there is a unique $Q^{*}$ function such that ${\mathcal{B}}^{\pi}_{r}[Q^{*}]=Q^{*}$, and for a $Q^{*}$ function in the $Q$-space, there is a unique reward function $r$ such that $r={\mathcal{T}}^{\pi}[Q^{*}]$. This result suggests that one can safely transform the objective function of the MaxEnt IRL from $r$-space to the Q-space as follows:
$\max_{Q}\min_{\pi}~{}~{}$ $\Phi(\pi,Q)=\mathbb{E}_{\rho}[{\mathcal{T}}^{\pi}[Q](s,a))]$
$-\mathbb{E}_{\rho_{\pi}}[{\mathcal{T}}^{\pi}[Q](s,a)]+\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]$ (3)
@@
which has several advantages; (Garg et al., 2021) show that $\Phi(\pi,Q)$ is convex in $\pi$ and linear in $Q$, implying that (3) always yields a unique saddle point solution. In particular, (3) can be converted into a maximization over the Q-space, making the training problem no longer adversarial.
3 Inverse soft-Q Learning with Suboptimal Demonstrations
We now describe our inverse soft-Q learning approach where we have not just expert trajectories, but also multiple levels of sub-optimal trajectories. First, we formulate an inverse soft-Q objective function that enables matching the occupancy distribution of expert and sub-optimal demonstrations through inverting a soft-Q function. Second, to further enhance the learning and mitigate the effect of limited expert samples, we introduce a regularization term in the inverse soft-Q objective. This additional term is to align the learning rewards with a pre-assigned reward function that allocates higher values to state-action pairs from higher expertise levels. Third, we show that this new objective does not have the same advantageous properties as the one in IQ-Learn (Garg et al., 2021), and we provide a modification to the new objective, which is a lower bound and has all the important properties.
3.1 Formulation of Inverse Soft-Q Objective
@@
We consider a setting where there are demonstrations classified into several sets of different expertise levels ${\mathcal{D}}^{1},{\mathcal{D}}^{2},....,{\mathcal{D}}^{N}$, where ${\mathcal{D}}^{1}$ consists of expert demonstrations and all the other sets contains sub-optimal ones. This setting is general and encompasses existing work in IL, which typically assumes that there are only two quality levels: expert and sub-optimal. Let ${\mathcal{D}}=\bigcup_{i\in[N]}{\mathcal{D}}^{i}$ be the union of all the demonstration sets and $\rho^{1},...,\rho^{N}$ be the occupancy measures of the respective expert policies. We then have an ordering of expected values across different levels of expert policies
@@
where $r^{*}(.,.)$ are the ground-truth rewards. Typically, we have the number of demonstrations in first level $\mathcal{D}^{1}$ is significantly lower than those from other expert levels, i.e.,
$|\mathcal{D}^{1}|\ll|\mathcal{D}^{i}|,\text{ for }i=2,...,N$
The MaxEnt IRL objective from Equation 2 can be adapted to account for the multiple expert level setting as
$\max_{r}\min_{\pi}~{}~{}$ $\sum_{i\in[N]}w_{i}\mathbb{E}_{\rho^{i}}[r(s,a)]-\mathbb{E}_{\rho_{\pi}}[r(s,a)]$
$\hskip 28.90755pt+\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]$ (4)
where $w_{i}\geq 0$ is the weight associated with the expert level $i\in[N]$ and we have $w_{1}>w_{2}>...>w_{N}$ and $\sum_{i\in[N]}w_{i}=1$. There are two key intuitions in the above optimization:
(a) Expert level $i$ accumulates higher expected values than expert levels greater than $i$;
(b) Difference in values accumulated by expert policies and the maximum of all other policies is maximized.
The optimization term can be rewritten as:
@@
where $\rho^{u}=\sum_{i\in[N]}w_{i}\rho^{i}$. Here we note that the expected reward $\sum_{i\in[N]}w_{i}\mathbb{E}_{\rho^{i}}[r(s,a)]$ is empirically approximated by samples from the demonstration sets ${\mathcal{D}}^{1},{\mathcal{D}}^{2},....,{\mathcal{D}}^{N}$, which may suffer from a low-sample size issue, as the number of samples in the best demonstration set $\mathcal{D}^{1}$ is low, so an empirical approximation of $\mathbb{E}_{\rho^{1}}[r(s,a)]$ using samples from ${\mathcal{D}}^{1}$ would be inaccurate. This is a foreseeable issue with any distributional matching approaches, especially when working with offline data without any online interactions. The primary challenge addressed in this section revolves around how to effectively utilize the optimal set of demonstrations $\mathcal{D}^{1}$ while mitigating the impact of low sample size.
3.2 Regularization with Reward Reference
To address the low-sample-size issue, we leverage the idea of the soft-Q Imitation Learning algorithm (Reddy et al., 2019). In this work, a reward reference of +1 is assigned to any state-action pair in the expert demonstrations, and the imitation is carried out by learning a policy that maximizes the expected accumulated pre-assigned rewards through soft-Q learning (so the learning policy will be trained towards replicating expert demonstrations). This approach would not suffer from a low-sample size issue and is more suitable to use with our inverse soft-Q approach, compared to other imitation learning algorithms that skip the reward inference step. To bring the idea into our context, we define a reward reference function $\overline{r}(s,a)$ such that $\overline{r}(s,a)>\overline{r}(s^{\prime},a^{\prime})$ for all $(s,a)$ in ${\mathcal{D}}^{1}$ and $(s^{\prime},a^{\prime})\notin{\mathcal{D}}^{1}$, and $\overline{r}(s,a)>\overline{r}(s^{\prime},a^{\prime})$ for all $(s,a)$ in ${\mathcal{D}}^{2}$ and $(s^{\prime},a^{\prime})\notin{\mathcal{D}}^{2}\cup{\mathcal{D}}^{1}$, and so on. The aim here is to assign higher rewards to higher expert demonstrations, and zero rewards to those that do not belong to any demonstrations. Keeping this in mind, we incorporate a reward regularization term into the MaxEntIRL objective in (4) as follows:
$\max_{r}\min_{\pi}\Big{\{}K(r,\pi)\stackrel{{\scriptstyle def}}{{=}}\underbrace{\sum_{i\in[N]}w_{i}\mathbb{E}_{\rho^{i}}[r(s,a)]-\mathbb{E}_{\rho_{\pi}}[r(s,a)]}_{\text{Occupancy matching}}$
$+\underbrace{\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]}_{\text{Entropy regularizer}}-\underbrace{\alpha\mathbb{E}_{\rho^{u}}[(r(s,a)-\overline{r}(s,a))^{2}]}_{\text{Reward regularizer}}\Big{\}}$ (5)

### v2

# SPRINQL: Sub-optimal Demonstrations driven Offline Imitation Learning

@@ Abstract
We focus on offline imitation learning (IL), which aims to mimic an expert’s behavior using demonstrations without any interaction with the environment. One of the main challenges in offline IL is the limited support of expert demonstrations, which typically cover only a small fraction of the state-action space. While it may not be feasible to obtain numerous expert demonstrations, it is often possible to gather a larger set of sub-optimal demonstrations. For example, in treatment optimization problems, there are varying levels of doctor treatments available for different chronic conditions. These range from treatment specialists and experienced general practitioners to less experienced general practitioners. Similarly, when robots are trained to imitate humans in routine tasks, they might learn from individuals with different levels of expertise and efficiency.
In this paper, we propose an offline IL approach that leverages the larger set of sub-optimal demonstrations while effectively mimicking expert trajectories. Existing offline IL methods based on behavior cloning or distribution matching often face issues such as overfitting to the limited set of expert demonstrations or inadvertently imitating sub-optimal trajectories from the larger dataset. Our approach, which is based on inverse soft-Q learning, learns from both expert and sub-optimal demonstrations. It assigns higher importance (through learned weights) to aligning with expert demonstrations and lower importance to aligning with sub-optimal ones. A key contribution of our approach, called SPRINQL, is transforming the offline IL problem into a convex optimization over the space of Q functions. Through comprehensive experimental evaluations, we demonstrate that the SPRINQL algorithm achieves state-of-the-art (SOTA) performance on offline IL benchmarks. Code is available at https://github.com/hmhuy2000/SPRINQL.
@@ 1 Introduction
Reinforcement learning (RL) has established itself as a strong and reliable framework for sequential decision-making with applications in diverse domains: robotics [19, 18], healthcare [34, 28, 27], and environment generation [9, 24]. Unfortunately, RL requires an underlying simulator that can provide rewards for different experiences, which is usually not available.
Imitation Learning (IL) [16, 29, 13, 17] handles the lack of reward function by utilizing expert demonstrations to guide the learning scheme to compute a good policy. However, IL approaches still require the presence of a simulator that allows for online interactions. Initial works in Offline IL  [35, 23, 40, 2] tackle the absence of simulator by considering an offline dataset of expert demonstrations. These approaches extend upon Behavioral Cloning (BC), where we aim to maximize the likelihood of the expert’s decisions from the provided dataset. The key advantage with BC is the theoretical justification on converging to expert behaviors given sufficient trajectories. However, when there are not enough expert trajectories, it often suffers from distributional shift issues [30]. Thus, a key drawback of these initial IL approaches is the need for a large number of expert demonstration datasets.
To deal with limited expert demonstrations, recent works utilize non-expert demonstration datasets to reduce the reliance on only expert demonstrations. These additional non-expert demonstrations are referred to as supplementary data. Directly applying BC to these larger supplementary datasets will lead to sub-optimal policies, so most prior work in utilizing supplementary data attempts to extract expert-like demonstrations from the supplementary dataset in order to expand the expert demonstrations [31, 21, 20, 37, 39]. These works assume expert-like demonstrations are present in the supplementary dataset and focus on identifying and utilizing those, while eliminating the non-expert demonstrations. Eliminating non-expert trajectories can result in loss of key information (e.g., transition dynamics) about the environment. Additionally, these works primarily rely on BC, which is known to overlook the sequential nature of decision-making problems – a small error can quickly accumulate when the learned policy deviates from the states experienced by the expert.
We develop our algorithm based on an inverse Q-learning framework that better captures the sequential nature of the decision-making [13] and can operate under the more realistic assumption that the data is collected from people/policies with lower expertise levels [^1] While there have been works [5, 6, 7] that have attempted to minimize the required demonstrations using ranked datasets by preference-based RL, their algorithms are only applicable to online settings. (not experts). To illustrate, consider a scenario in robotic manipulation where the goal is to teach a robot to assemble parts. Expert demonstrations might show precise and efficient methods to assemble parts, but are limited in number due to the high cost and time associated with expert involvement. On the other hand, sub-optimal demonstrations from novice users are easier to obtain and more abundant. Our SPRINQL approach effectively integrates these sub-optimal demonstrations, giving appropriate weight to the expert demonstrations to ensure the robot learns the optimal assembly method without overfitting to the limited expert data or the inaccuracies in the sub-optimal data. We utilize these non-expert trajectories to learn a Q function that contributes to our understanding of the environment and the ground truth reward function.
Contributions: Overall, we make the following key contributions in this paper:
@@
(i) We propose SPRINQL, a novel algorithm based on Q-learning for offline imitation learning with expert and multiple levels of sub-optimal demonstrations.
(ii) We provide key theoretical properties of the SPRINQL objective function, which enable the development of a scalable and efficient approach. In particular, we leverage distribution matching and reward regularization to develop an objective function for SPRINQL that not only help address the issue of limited expert samples but also utilizes non-expert data to enhance learning. Our objective function is not only convex within the space of $Q$ functions but also guarantees the return of a $Q$ function that lower-bounds its true value.
(iii) We provide an extensive empirical evaluation of our approach in comparison to existing best algorithms for offline IL with sup-optimal demonstrations. Our algorithms provide state-of-the-art (SOTA) performance on all the benchmark problems. Moreover, SPRINQL is able to recover a reward function that shows a high positive correlation with the ground-truth rewards, highlighting a unique advantage of our approach compared to other IL algorithms in this context.
@@ 1.1 Related Work
Imitation learning is recognized as a significant technique for learning from demonstrations. It begins with BC, which aims to maximize the likelihood of expert demonstrations. However, BC often under-performs in practice due to unforeseen scenarios [30]. To overcome this limitation, Generative Adversarial Imitation Learning (GAIL) [16] and Adversarial Inverse Reinforcement Learning (AIRL) [10] have been developed. These methods align the occupancy distributions of the policy and the expert within the Generative Adversarial Network (GAN) framework [14]. Alternatively, Soft Q Imitation Learning (SQIL) [29] bypasses the complexities of adversarial training by assigning a reward of +1 to expert demonstrations and 0 to the others, subsequently learning a value function based on these rewards.While the aforementioned imitation learning algorithms show promise, they require interaction with the environment to obtain the policy distribution, which is often impractical.
@@
ValueDICE [22] introduces a novel approach for off-policy training, suitable for offline training, using Stationary Distribution Corrections [25, 26]. However, ValueDICE necessitates adversarial training between the policy network and the $Q$ network, which can make the training slow and unstable. Recently, algorithms like PWIL [8] and IQ-learn [13] have optimized distribution distance, offering an alternative to adversarial training schemes. Since such approaches rely on occupancy distribution matching, a large expert dataset is often required to achieve the desired performance. Our approach, SPRINQL is able to bypass this requirement of a large set of expert demonstrations through the use of non-expert demonstrations (which are typically more available) in conjunction with a small set of expert demonstrations.
Imitation Learning with Imperfect Demonstrations.
T-REX [5] and D-REX [6] have shown that utilizing noise-ranked demonstrations as a reference-based approach can return a better policy without requiring expert demonstrations in online settings. Moreover, there are also several works [36, 33] that utilize the GAN framework [14] for sub-optimal datasets and have achieved several successes. Meanwhile, in the offline imitation learning context, TRAIL [38] utilizes sup-optimal demonstrations to learn the environment’s dynamics. It employs a feature encoder to map the high-dimensional state-action space into a lower dimension, thereby allowing for a scalable way of learning of dynamics. This approach may face challenges in complex environments where predicting dynamics accurately is difficult, as shown in our experimental results. Other works assume that they can extract expert-like state-action pairs from the sub-optimal demonstration set and use them for BC with importance sampling [31, 37, 39, 21, 20]. However, expert-like state-actions might be difficult to accurately identify, as true reward information is not available. In contrast, our approach is more general, as we do not assume that the sub-optimal set contains expert-like demonstrations. We also allow for the inclusion of demonstrations of various qualities. Moreover, while prior works only recover policies, our approach enables the recovery of both expert policies and rewards, justifying the use of our method for Inverse Reinforcement Learning (IRL)[1].
@@ 2 Background
We consider a MDP defined by the following tuple $\mathcal{M}=\left\langle S,A,r,P,\gamma,s_{0}\right\rangle$, where $S$ denotes the set of states, $s_{0}$ represents the initial state set, $A$ is the set of actions, $r:S\times A\rightarrow\mathbb{R}$ defines the reward function for each state-action pair, and $P:S\times A\rightarrow S$ is the transition function, i.e., $P(s^{\prime}|s,a)$ is the probability of reaching state $s^{\prime}\in S$ when action $a\in A$ is made at state $s\in S$, and $\gamma$ is the discount factor. In reinforcement learning (RL), the aim is to find a policy that maximizes the expected long-term accumulated reward $\max_{\pi}~\left\{\mathbb{E}_{(s,a)\sim\rho_{\pi}}[r(s,a)]\right\}$, where $\rho_{\pi}$ is the occupancy measure of policy $\pi$: $\rho_{\pi}(s,a)=(1-\gamma)\pi(a|s)\sum_{t=1}^{\infty}\gamma^{t}P(s_{t}=s|\pi).$
@@
The objective in MaxEnt IRL is to recover a reward function $r(s,a)$ from a set of expert demonstrations, ${\mathcal{D}}^{E}$. Let $\rho^{E}$ be the occupancy measure of the expert policy. The MaxEnt IRL framework [41] proposes to recover the expert reward function by solving
@@
Intuitively, the aim is to find a reward function that achieves the highest difference between the expected value of the expert policy and the highest expected value among all other policies (computed through the min loop).
@@
Given a reward function $r$ and a policy $\pi$, the soft Bellman equation is defined as ${\mathcal{B}}^{\pi}_{r}[Q](s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}}[V^{\pi}(s^{\prime})],$ where $V^{\pi}(s)=\mathbb{E}_{a\sim\pi(a|s)}[Q(s,a)-\log\pi(a|s)]$. The Bellman equation ${\mathcal{B}}^{\pi}_{r}[Q]=Q$ is contractive and always yields a unique Q solution [13]. In IQ-learn, they further define an inverse soft-Q Bellman operator ${\mathcal{T}}^{\pi}[Q]=Q(s,a)-\gamma\mathbb{E}_{s^{\prime}}[V^{\pi}(s^{\prime})]$. [13] show that for any reward function $r(a,s)$, there is a unique $Q^{*}$ function such that ${\mathcal{B}}^{\pi}_{r}[Q^{*}]=Q^{*}$, and for a $Q^{*}$ function in the $Q$-space, there is a unique reward function $r$ such that $r={\mathcal{T}}^{\pi}[Q^{*}]$. This result suggests that one can safely transform the objective function of the MaxEnt IRL from $r$-space to the Q-space as follows:
$\max_{Q}\min_{\pi}~{}~{}$ $\Phi(\pi,Q)=\mathbb{E}_{\rho}[{\mathcal{T}}^{\pi}[Q](s,a))]-\mathbb{E}_{\rho_{\pi}}[{\mathcal{T}}^{\pi}[Q](s,a)]+\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]$ (2)
@@
which has several advantages; [13] show that $\Phi(\pi,Q)$ is convex in $\pi$ and linear in $Q$, implying that (2) always yields a unique saddle point solution. In particular, (2) can be converted into a maximization over the Q-space, making the training problem no longer adversarial.
3 SPRINQL
We now describe our inverse soft-Q learning approach, referred to as SPRINQL (Sub-oPtimal demonstrations driven Reward regularized INverse soft Q Learning). We first describe the three key components in the SPRINQL formulation:
(1) We formulate the objective function that enables matching the occupancy distribution of not just expert demonstrations, but also sub-optimal demonstrations.
(2) To mitigate the effect of limited expert samples (and larger sets of sub-optimal samples) that can bias the distribution matching of the first step to sub-optimal demonstrations, we introduce a reward regularization term within the objective. This regularization term is to ensure reward function allocates higher values to state-action pairs that appear in higher expertise demonstrations.
(3) We show that while this new objective does not have the same advantageous properties as the one in inverse Q-learning [13], with some minor (yet significant) changes it is possible to restore all the important properties.
3.1 Distribution Matching with Expert and Suboptimal Demonstrations
@@
We consider a setting where there are demonstrations classified into several sets of different expertise levels ${\mathcal{D}}^{1},{\mathcal{D}}^{2},....,{\mathcal{D}}^{N}$, where ${\mathcal{D}}^{1}$ consists of expert demonstrations and all the other sets contains sub-optimal ones. This setting is general than existing work in IL with sup-optimal demonstrations, which typically assumes that there are only two quality levels: expert and sub-optimal. Let ${\mathcal{D}}=\bigcup_{i\in[N]}{\mathcal{D}}^{i}$ be the union of all the demonstration sets and $\rho^{1},...,\rho^{N}$ be the occupancy measures of the respective expert policies. The ordering of expected values across different levels of expert policies would then be given by:
@@
where $r^{*}(.,.)$ are the ground-truth rewards. Typically, the number of demonstrations in first level, $\mathcal{D}^{1}$ is significantly lower than those from other expert levels, i.e., $|\mathcal{D}^{1}|\ll|\mathcal{D}^{i}|,\text{ for }i=2,...,N$ The MaxEnt IRL objective from Equation 1 can thus be adapted as follows:
$\max_{r}\min_{\pi}~{}~{}$ $\sum_{i\in[N]}w_{i}\mathbb{E}_{\rho^{i}}[r(s,a)]-\mathbb{E}_{\rho_{\pi}}[r(s,a)]+\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]$ (3)
where $w_{i}\geq 0$ is the weight associated with the expert level $i\in[N]$ and we have $w_{1}>w_{2}>...>w_{N}$ and $\sum_{i\in[N]}w_{i}=1$. There are two key intuitions in the above optimization: (a) Expert level $i$ accumulates higher expected values than expert levels greater than $i$; and (b) Difference in values accumulated by expert policies and the maximum of all other policies is maximized. The optimization term can be rewritten as:
@@
where $\rho^{U}=\sum_{i\in[N]}w_{i}\rho^{i}$. Here we note that the expected reward $\sum_{i\in[N]}w_{i}\mathbb{E}_{\rho^{i}}[r(s,a)]$ is empirically approximated by samples from the demonstration sets ${\mathcal{D}}^{1},{\mathcal{D}}^{2},....,{\mathcal{D}}^{N}$. The number of demonstrations in the best demonstration set $\mathcal{D}^{1}$ (i.e. the set of expert demonstrations) is significantly lower when compared to other demonstration sets. So, an empirical approximation of $\mathbb{E}_{\rho^{1}}[r(s,a)]$ using samples from ${\mathcal{D}}^{1}$ would be inaccurate.
3.2 Regularization with Reference Reward
We create a reference reward based on the provided expert and sub-optimal demonstrations and utilize the reference function to compute a regularization term that is added to the objective of Equation 3. Concretely, we define a reference reward function $\overline{r}(s,a)$ such that:
$\overline{r}(s,a)>\overline{r}(s^{\prime},a^{\prime}),\forall(s,a)\in{\mathcal{D}}^{1}\text{ and }(s^{\prime},a^{\prime})\notin{\mathcal{D}}^{1}\text{ and }$
$\overline{r}(s,a)>\overline{r}(s^{\prime},a^{\prime}),\forall(s,a)\in{\mathcal{D}}^{2}\text{ and }\forall(s^{\prime},a^{\prime})\notin{\mathcal{D}}^{2}\cup{\mathcal{D}}^{1}\text{ and so on}$
The aim here is to assign higher rewards to demonstrations from higher expertise levels, and zero rewards to those that do not belong to provided demonstrations. We will discuss how to concretely estimate such reference reward values later.
We utilize this reference reward as part of the reward regularization term, which is added into the MaxEntIRL objective in (3) as follows:
[…]